OpenStudy (anonymous):

Given that cosΘsinΘ=1/2, evaluate (cosΘ-sinΘ)^2

OpenStudy (anonymous):

this is like in another language ! ugh i hate math -__-

OpenStudy (anonymous):

Lol yeah. Do you understand it though?

OpenStudy (raden):

let the value of (cosΘ-sinΘ)^2 = t, just expand it for left side gives cos^2 Θ + sin^2 Θ - 2sinΘcosΘ = t 1 - 2*1/2 = t 1-1=t so, t=0

OpenStudy (anonymous):

Wait, so where did the 1 - 2*1/2 = t come from?

OpenStudy (raden):

u must know that identity of cos^2 Θ + sin^2 Θ = 1 and the value of sin Θcos Θ is given :)

OpenStudy (anonymous):

1) he multiplied it out 2) trig identities: sin^u+cost^2u = 1 3) 1/2 comes from the quiestion itself

OpenStudy (anonymous):

OOOHH TRIG IDENTITIES i FORGOT ABOUT THOSE!

OpenStudy (anonymous):

that should be sin^2u

OpenStudy (raden):

nopes, @ivan77 ... i see :) study hard again @xxxtechmanxxx :p

OpenStudy (anonymous):

Lawls, yeah, Pre-Calc isnt my strognest subject :P

OpenStudy (raden):

hehe... no problems, let's enjoying study in openstudy. because it's free :)

OpenStudy (anonymous):

openstudy lags so badly for me tht I cant stand it anymore.